OpenStudy (anonymous):

how many ways are there of choosing 2 letters, without replacement, from the 5 letters A, B, C, D, and E, if the order of the choices matters? the order of the choices does not matter?

OpenStudy (mathstudent55):

In these types of problems, if the order matters, it's a permutation problem. If the order does not matter, it's a combination problem.

OpenStudy (anonymous):

well it si asking for both..... I dont understand how to do it and my math teacher did not teach this at all in class..... so 

OpenStudy (mathstudent55):

In the permutation problem, you need to choose all permutations of 2 letters from 5. That is \(\large_5P_2\). The formula for permutation is: \(\large _nP_r = \dfrac{n!}{(n - r)!} \)

OpenStudy (anonymous):

WHAT :#

OpenStudy (mathstudent55):

You are choosing all permutations of 2 items chosen from 5. In your case, n = 5, and r = 2. Just put them in the formula above and evaluate it. Are you familiar with factorial numbers?

OpenStudy (anonymous):

NNOOOO

OpenStudy (anonymous):

!!!!!

OpenStudy (mathstudent55):

Ok. Then let me explain that to you. A factorial number is this: 1! (read as one factorial), 2! (read as 2 factorial) etc. The definition is 1! = 1 2! = 2 * 1 = 2 3! = 3 * 2 * 1 = 6 4! = 4 * 3 * 2 * 1 = 24 etc.

OpenStudy (anonymous):

oh... so 7!: 1,7 and 6!: 1,2,3,6

OpenStudy (mathstudent55):

An integer greater than 1 factorial is equal to that integer multiplied by all integers less than that number and greater than or equal to one.

OpenStudy (anonymous):

or is it.....

OpenStudy (mathstudent55):

No. Let's do one thing at a time. For now we are just learning about factorial. Then we will go over the formula for permutations.

OpenStudy (mathstudent55):

7! = 7 * 6 * 5 * 4 * 3 * 2 * 1 = 5040

OpenStudy (anonymous):

so.................... 7!: 7*6*5*4*3*2*1 and 5!: 5*4*3*2*1

OpenStudy (mathstudent55):

6! = 6 * 5 * 4 * 3 * 2 * 1 = 720

OpenStudy (mathstudent55):

Correct. Now you got it!

OpenStudy (anonymous):

YAY

OpenStudy (anonymous):

so

OpenStudy (mathstudent55):

Ok. That is how factorial numbers work. Now we can go on to the formula for the number of permutations. If you are choosing 2 items out of 5, where order matters, the formula is: \(\large _nP_r = \dfrac{n!}{(n - r)!}\) for choosing r items out of n items where order matters.

OpenStudy (mathstudent55):

In your case, n = 5, and r = 2, so we can plug in those numbers in the formula to get: \(\large _5P_2 = \dfrac{5!}{(5 - 2)!}\) Now we need to simplify the formula until we get a number.

OpenStudy (mathstudent55):

\(\large _5P_2 = \dfrac{5!}{(5 - 2)!} = \dfrac{5!}{3!} = \dfrac{5 \times 4 \times 3 \times 2 \times 1}{3 \times 2 \times 1}\)

OpenStudy (mathstudent55):

Ok so far?

OpenStudy (anonymous):

yeah.

OpenStudy (mathstudent55):

We need a number at the end, not an operation, so we need to continue simplifying.

OpenStudy (anonymous):

so it's 120/6

OpenStudy (anonymous):

20

OpenStudy (mathstudent55):

\(=\dfrac{5 \times 4 \times \cancel{3} \times \cancel{2} \times \cancel{1}}{\cancel{3} \times \cancel{2} \times \cancel{1}} = 5 \times 4 = 20\)

OpenStudy (mathstudent55):

Correct.

OpenStudy (mathstudent55):

Now we need to do the second part. If order does not matter, then we will get fewer results. Then we are dealing with a combination problem. The formula for the number of combibnations of choosing r items from n items is: \(_nC_r = \dfrac{n!}{(n - r)!r!} \)

OpenStudy (mathstudent55):

Now try this formula with n = 5 and r = 2, and see what you get.

OpenStudy (anonymous):

@mathstudent55 I am sorry I had to eat dinner ;) lol

OpenStudy (mathstudent55):

No problem. Just try the combination formula with n = 5 and r = 2, and let me know what you get.